# Solving and learning advective multiscale Darcian dynamics with the Neural Basis Method

### v1

@@ Abstract
Physics-governed models are increasingly paired with machine learning for accelerated predictions, yet most “physics–informed” formulations treat the governing equations as a penalty loss whose scale and meaning are set by heuristic balancing. This blurs operator structure, thereby confounding solution approximation error with governing-equation enforcement error and making the solving and learning progress hard to interpret and control. Here we introduce the Neural Basis Method, a projection-based formulation that couples a predefined, physics-conforming neural basis space with an operator-induced residual metric to obtain a well-conditioned deterministic minimization. Stability and reliability then hinge on this metric: the residual is not merely an optimization objective but a computable certificate tied to approximation and enforcement, remaining stable under basis enrichment and yielding reduced coordinates that are learnable across parametric instances. We use advective multiscale Darcian dynamics as a concrete demonstration of this broader point. Our method produce accurate and robust solutions in single solves and enable fast and effective parametric inference with operator learning.
@@
Keywords: neural basis method, operator learning, residual-error equivalence, parametric PDEs, coupled Darcy flow and transport, multiscale advection-diffusion problem
footnotetext: ∗†These authors contribute equally to this work.
Introduction
@@
Flow–transport coupling in multiscale porous materials [35, 21] pose major challenges for computational modeling, especially in parametric many-query settings [1, 32]. Governed by elliptic–parabolic and advective PDEs with heterogeneous coefficients, such coupled dynamics entail conservation constraints, nonlinear effects, and multiscale interactions [29, 45]. Physics-informed machine learning [30], such as physics-informed neural networks (PINNs) [41, 47], is often presented as promising because it offers mesh-free approximations, seamless data integration, and potential learning-enabled acceleration. In practice, however, most approaches enforce physics through dimensionless $L^{2}$ loss minimization with heuristic weighting, which obscures physical scaling and disrupts numerical structure. This can yield poorly conditioned optimization in which approximation error is entangled with operator inconsistency, undermining accuracy and reliability [48, 5, 33, 51, 50]. This raises a fundamental question: where should physics and numerical structure reside in learning-based PDE models?
@@
In this work, we propose the Neural Basis Method (NBM), which departs from loss-driven training by seeking solutions through an explicit, deterministic procedure. We use predefined physics-conforming neural networks to construct finite-dimensional approximation spaces. Then, we impose PDEs by formulating a well-defined projection problem and solve for the solution coefficients. This design aligns NBM with classical numerical analysis and allows us to incorporate variational formulations [52] in a systematic and principled manner. Furthermore, for parametric cases, we represent solution families across parameter instances within the same neural space by learning the parameter-dependent solution coefficients, which naturally yields an operator learning formulation.
@@
NBM is motivated by two lines of ideas: projection-based methods (e.g., Galerkin), which use fixed finite-dimensional space to enable stability, error control, and systematic refinement [8, 2], and neural approximation approaches, which offer expressive spaces for representing solution manifolds [15, 3, 53]. NBM therefore aims to integrate neural representations while retaining the structure and guarantees of projection-based schemes. Related ideas have recently appeared, including random feature and extreme learning machine-type methods, where fixed single-hidden-layer neural networks serve as basis functions and solution coefficients are obtained by least-squares [28, 11, 40, 20, 57]. They provide an important proof-of-concept, but the non-orthogonal bases often yield severely ill-conditioned least-squares systems [13, 54, 44, 12]. Especially, for multiscale problems, resolving fine-scale structures requires basis enrichment [11, 27, 55, 14], which can drive rapid, often exponential, condition-number growth. NBM builds on this direction while departing in three ways: a multilayer neural basis generator that resists conditioning degradation under basis enrichment; an essential physical structure that is built in at vector field representation level; and a weighting and stabilization strategy that enforces consistency with the underlying operator and physical scaling. Our design draws on insights from least-squares finite element methods (LSFEM) and the discontinuous Petro-Galerkin (DPG) framework. LSFEM inspires properly scaled boundary residuals in seeking solutions via operator-residual minimization [6, 43, 9, 10]. DPG theory connects optimal test norms for multiscale and advective problems to generalized least-squares formulations [19, 17]. Collectively, these insights motivate the NBM view that weighting and stabilization are treated as principled components of the projection-based approach rather than as heuristic loss balancing, preserving operator consistency and physical scaling and enabling reliable residual-based measures.
@@
We then extend NBM from single instance solves to parametric operator learning (NBM-OL). In many query settings, including boundary/source control, performance optimization, and inverse modeling, the goal is to learn how the solution operator varies with governing parameters [34, 49, 31]. This necessitates a measure that is numerically comparable across parameters and, under stable projection, provides a computable a posteriori bound on the solution error [25, 38]. In NBM-OL, we learn parametric dependence in the fixed neural basis space equipped with such measure, instead of surrogate losses that lack direct numerical meaning. This provides a robust utility for monitoring training and steering refinement across hyperparameters, with well-informed stopping criterion.
We present the first complete formulation and demonstration of NBM and NBM-OL in the setting of advective multiscale Darcian dynamics. For this coupled Darcy flow–transport system, we use physics-conforming neural vector bases in a nonlinear mixed Darcy flow formulation with energy-consistent weighting, and advance transport using a purely advective update stabilized by upwind control-volume treatment. Numerical experiments show accurate and stable single-instance solutions, and, for parametric inference, excellent in-distribution generalization, strong out-of-distribution robustness, and massive acceleration. In representative examples, we observe speedups on the order of $10^{3}$$10^{4}$, and the gain scales with the prediction temporal horizon and problem size. Although we use a collocation-based realization for clarity and efficiency, our method also supports variational formulations, including DPG-inspired constructions with the optimal test theory that can provide theoretically optimal weighting and stabilization for multiscale and advective problems, with potential adaptive error control via the Riesz representation theorem [16]. A dedicated, DPG-inspired development of NBM is reported in subsequent, specialized work.
Results
We present the key progress of NBM development through an instantiation of coupled Darcy flow–transport in multiscale porous materials as a concrete testbed (descriptions in Methods). Using this context, we first summarize the key ingredients of NBM for solving the system; see Fig. 1. We next demonstrate NBM’s behavior on single-query setups. We then consider parametric many-query regimes, where variability arises from permeability fields and boundary values, and show how parametric inference is achieved through NBM-OL.
@@ Results
Overview of NBM as a neural PDE solver
Core NBM formulation. NBM approximates solutions with finite neural basis functions and computes the solution coefficients through a PDE-constrained least-squares projection. For system (M1)-(M2), the pressure $p(\mathbf{x},t)$ and mass-flux $\mathbf{q}(\mathbf{x},t)$ are smooth and therefore suit global basis expansion naturally. In contrast, the concentration $c(\mathbf{x},t)$ can develop limited regularity with discontinuous fronts and thus requires specialized stabilization or regularization. In general, we write $s_{NN}(\mathbf{x},t)=\sum\theta_{i}(t)\,\phi_{i}(\mathbf{x})$, where a multilayer, fixed-parameter, ResNet-style network [26] generates the scalar spatial bases $\{\phi_{i}(\mathbf{x})\}$. Moreover, we let the coefficient vector $\boldsymbol{\theta}(t)$ carry temporal evolution. Here $s$ denotes a scalar field, namely pressure, concentration, or a mass-flux component such as $q_{x}$ or $q_{y}$. We freeze network weights/biases, so we defines a deterministic approximation space $\mathrm{span}\{\phi_{i}\}$ and obtain the solution by projecting onto this space rather than by training the network. Specifically, we use two hidden layers to balance expressiveness and numerical conditioning while maintaining a clear geometric interpretation of the induced basis functions. We refer to this construction as a dual-layer neural basis; see Methods and SI.2-5 for details. For the vector field $\mathbf{q}$, we propose a neural vector basis design inspired by the Helmholtz decomposition [42, 46] such that we approximate it as a sum of a divergence-free part $\boldsymbol{\phi}_{i}^{(\mathrm{div})}$ and curl-free part $\boldsymbol{\phi}_{i}^{(\mathrm{div})}$,

### v2

@@ Abstract
Physics–governed models are increasingly paired with machine learning for accelerated computations, yet most “physics–informed” formulations treat the governing equations as a penalty loss whose scale and meaning are set by heuristic balancing. This blurs operator structure, thereby confounding solution approximation error with governing-equation enforcement error and making the solving and learning progress hard to interpret and control. Here we introduce the Neural Basis Method, a projection-based formulation that couples a predefined, physics-conforming neural basis space with an operator-induced residual metric to obtain a well-conditioned deterministic minimization. Stability and reliability then hinge on this metric: the residual is not merely an optimization objective but a computable certificate tied to approximation and enforcement, remaining stable under basis enrichment and yielding reduced coordinates that are learnable across parametric instances. We use advective multiscale Darcian dynamics as an engineering demonstration to illustrate this broader point. Our method produces accurate and robust solutions in single solves and enables fast and effective parametric inference for essential engineering tasks such as control, optimization and uncertainty quantification.
@@
Keywords: neural basis method, operator learning, parametric inference, reduced-order modeling, Darcy flow and transport, multiscale flow simulation
1. Introduction
@@
Flow–transport coupling in multiscale porous materials [35, 21] poses major challenges for modeling, especially in parametric many-query workflows [1, 32]. Governed by elliptic–parabolic and advective PDEs with heterogeneous coefficients, such multiphysics dynamics entail conservation constraints, nonlinear effects, and multiscale interactions [29, 48]. Physics-informed machine learning [30], such as physics-informed neural networks (PINNs) [41, 50], is often presented as promising because it offers mesh-free approximations, seamless data integration, and potential learning-enabled acceleration. In practice, however, most approaches enforce physics through dimensionless $L^{2}$ loss minimization with heuristic weighting, which obscures physical scaling and disrupts numerical structure. This can yield poorly conditioned optimization in which approximation error is entangled with operator inconsistency, undermining accuracy and reliability [51, 5, 33, 54, 53]. This raises a fundamental question: where should physics and numerical structure reside in learning-based PDE models?
@@
In this work, we propose the Neural Basis Method (NBM), which departs from loss-driven training by seeking solutions through an explicit, deterministic procedure. We use predefined physics-conforming neural networks to construct finite-dimensional approximation spaces. Then, we impose PDEs by formulating a well-defined projection problem and solve for the solution coefficients. This design aligns NBM with classical numerical analysis and allows us to incorporate variational formulations [55] in a systematic and principled manner. Furthermore, for parametric cases, we represent solution families across parameter instances within the same neural space by learning the parameter-dependent solution coefficients, which naturally yields an operator learning formulation. This perspective is especially useful in many-query engineering settings under varying coefficients, boundary conditions, or control inputs.
@@
NBM is motivated by two lines of ideas: projection-based methods (e.g., Galerkin), which use fixed finite-dimensional space to enable stability, error control, and systematic refinement [8, 2], and neural approximation approaches, which offer expressive spaces for representing solution manifolds [15, 3, 56]. NBM therefore aims to integrate neural representations while retaining the structure and guarantees of projection-based schemes. Related ideas have recently appeared, including random feature and extreme learning machine-type methods, where fixed single-hidden-layer neural networks serve as basis functions and solution coefficients are obtained by least-squares [28, 11, 40, 20, 60, 43, 44, 46]. They provide an important proof-of-concept, but the non-orthogonal bases often yield severely ill-conditioned least-squares systems [13, 57, 47, 12]. Especially, for multiscale problems, resolving fine-scale structures requires basis enrichment [11, 27, 58, 14], which can drive rapid, often exponential, condition-number growth. NBM builds on this direction while departing in three ways: a multilayer neural basis generator that resists conditioning degradation under basis enrichment; an essential physical structure that is built in at vector field representation level; and a weighting and stabilization strategy that enforces consistency with the underlying operator and physical scaling. Our design draws on insights from least-squares finite element methods (LSFEM) and the discontinuous Petro-Galerkin (DPG) framework. LSFEM inspires properly scaled boundary residuals in seeking solutions via operator-residual minimization [6, 45, 9, 10]. DPG theory connects optimal test norms for multiscale and advective problems to generalized least-squares formulations [19, 17]. Collectively, these insights motivate the NBM view that weighting and stabilization are treated as principled components of the projection-based approach rather than as heuristic loss balancing, preserving operator consistency and physical scaling and enabling reliable residual-based measures.
@@
We then extend NBM from single instance solves to parametric operator learning (NBM-OL). In many-query engineering tasks, including control, optimization, and uncertainty quantification, the goal is to learn how the solution operator varies with governing parameters [34, 52, 31]. This necessitates a measure that is numerically comparable across parameters and, under stable projection, provides a computable a posteriori bound on the solution error [25, 38]. In NBM-OL, we learn parametric dependence in the fixed neural basis space equipped with such measure, instead of surrogate losses that lack direct numerical meaning. This provides a robust utility for monitoring training and steering refinement across hyperparameters, with well-informed stopping criterion.
In this study, we present the first complete formulation and demonstration of NBM/NBM-OL in the setting of advective multiscale Darcian dynamics, with carbon storage as a representative application. We use physics-conforming neural vector bases in a nonlinear mixed Darcy flow formulation with energy-consistent weighting, and advance transport using a purely advective update stabilized by upwind control-volume treatment. As an introductory paper, we adopt a collocation-based realization for clarity and efficiency, although the method also supports variational formulations. Numerical experiments show accurate and stable single-instance solutions, and, for parametric inference, excellent in-distribution generalization, strong out-of-distribution robustness, and massive acceleration. In representative examples, we observe computational speedups on the order of $10^{3}$$10^{4}$, and the gain scales with the prediction temporal horizon and problem size.
This paper is organized as follows. In the setting of coupled Darcy flow–transport, we first present the core ingredients of NBM and evaluate its performance on single-instance problems. We then turn to parametric many-query regimes, where variability arises from permeability fields and boundary conditions, and show how NBM-OL enables efficient parametric inference. Together, we illustrate the central principle of NBM: by placing physics in the neural discretization through physics-conforming approximation spaces and operator-induced residual metrics, the method unifies direct PDE solving and parametric operator learning within the same projection-based formulation.
2. Overview of NBM as a neural PDE solver
The overall workflow of NBM for coupled multiscale Darcy flow–transport (see Appendix A) is summarized in Fig. 1. It brings together the main components of the collocation-based implementation, including the underlying physical system, neural basis approximations for scalar (pressure, concentration) and vector fields (mass-flux), the mixed weighted least-squares formulation for the multiscale Darcy flow solution with energy-consistent weighting of interior and boundary residuals, and the implicit time-marching procedure for the coupled flow–transport dynamics.
@@ 2. Overview of NBM as a neural PDE solver
2.1 Core NBM formulation
NBM approximates solutions with finite neural basis functions and computes the solution coefficients through a PDE-constrained least-squares projection. For system (A.1)-(A.2), the pressure $p(\mathbf{x},t)$ and mass-flux $\mathbf{q}(\mathbf{x},t)$ are smooth and therefore suit global basis expansion naturally. In contrast, the concentration $c(\mathbf{x},t)$ can develop limited regularity with discontinuous fronts and thus requires specialized stabilization or regularization. In general, we write $s_{NN}(\mathbf{x},t)=\sum\theta_{i}(t)\,\phi_{i}(\mathbf{x})$, where a multilayer, fixed-parameter, ResNet-style network [26] generates the scalar spatial bases $\{\phi_{i}(\mathbf{x})\}$. Moreover, we let the coefficient vector $\boldsymbol{\theta}(t)$ carry temporal evolution. Here $s$ denotes a scalar field, namely pressure, concentration, or a mass-flux component such as $q_{x}$ or $q_{y}$. We freeze network weights/biases, so we defines a deterministic approximation space $\mathrm{span}\{\phi_{i}\}$ and obtain the solution by projecting onto this space rather than by training the network. Specifically, we use two hidden layers to balance expressiveness and numerical conditioning while maintaining a clear geometric interpretation of the induced basis functions. We refer to this construction as a dual-layer neural basis.
For the vector field $\mathbf{q}$, we propose a neural basis design inspired by the Helmholtz decomposition [42, 49] such that we approximate via a sum of a divergence-free part $\boldsymbol{\phi}_{i}^{(\mathrm{div})}$ and curl-free part $\boldsymbol{\phi}_{i}^{(\mathrm{div})}$,
@@ 2.1 Core NBM formulation
We build the curl-free part as $\boldsymbol{\phi}_{i}^{(\mathrm{curl})}=\nabla\phi_{i}$, which represents the irrotational flow driven by potential. For the divergence-free part, we use a streamfunction formulation in 2D, $\boldsymbol{\phi}_{i}^{(\mathrm{div})}=\nabla^{\perp}\phi_{i}$, and a vector potential formulation in 3D, $\boldsymbol{\phi}_{i}^{(\mathrm{div})}=\nabla\times\mathbf{F}_{i}$, where $\mathbf{F}_{i}$ can be constructed from the same scalar basis $\{\phi_{i}\}$. By this design, $\nabla\cdot\boldsymbol{\phi}_{i}^{(\mathrm{div})}=0$ holds identically, so only the curl-free subspace enters equation (A.2); it supplies the degrees of freedom needed to represent compressible effects and source-driven mass redistribution, while the divergence-free component naturally captures mass-preserving motion. This physics-conforming representation preserves the intrinsic structure of fluid motion as much as possible at the representation level and can effectively improve solution accuracy. See Appendices B-E for details of construction and properties of neural basis functions.